# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.1.5 Results
In general, the authors of the papers in this group claim that their proposed GNN model outperforms the benchmarks. However, in some cases, it turns out that basic statistical methods may achieve good results as well [art_390]. The most common error metrics used are the MAE, the RMSE, and the Normalized Root mean squared Error (NRMSE). Some papers calculate these error metrics across different seasons to evaluate the model’s performance under different conditions. An example is provided in Ref. [art_349], where the authors also use the GNNExplainer [interpret_2] to assess the influence of different variables.
It is not possible to compare the accuracy of the models included in this group, because no common benchmark datasets were used. The only two papers that study the same dataset are written by the same authors (Refs. [art_21, art_22]), and their most recent paper includes a comparison with the models proposed in the first one.
@@ 5.2 Environment
The thematic group “Environment” includes a relevant number of selected papers (43 out of 281). This suggests a considerable research interest on the application of GNNs in this field. However, this group is highly fragmented, as it includes a large number of subdomains and applications, all related to the study of environmental data.
@@ 5.2.1 Overview
The group of papers related to environmental topics is broad, and includes papers on disciplines such as physics, climatology, oceanography, and atmospheric science. The breadth of the group is reflected in the number of case studies, which cover air quality prediction (in terms of PM2.5 concentration), sea temperature, wind speed, and other applications. Forecasting these quantities can be useful for several reasons. For instance, predicting wind speed or rainfall can assist in anticipating some extreme weather phenomena, and is also related to the prediction of green energy outputs. For example, in Refs. [art_29] and [art_34] the authors estimate the wind power forecasting error obtained by using wind speed forecasting models. As another example, predicting sea surface temperature can be useful for weather forecasting, fishing directions, and disaster warnings [art_28]. Moreover, monitoring and forecasting environmental data are crucial for evaluating the impact of human activities on the environment and tracking the progression of climate change.
@@
An overview of the selected papers reveals that there has been a rapid increase in interest on this theme among the GNN research community too, and more than 35 of the selected papers were published from 2023. The most investigated data are related to sea temperature (Refs. [art_28, art_33, art_35, art_36, art_39, art_41, art_47]), wind speed (Refs. [art_29, art_31, art_34, art_40, art_44, art_46, art_366]), and PM2.5 concentration forecasting (Refs. [art_2, art_3, art_4, art_6, art_7, art_8, art_354, art_361, art_371]). Almost every selected paper is published in a different journal, with the exception of Energy, Engineering Applications of Artificial Intelligence, Journal of Cleaner Production, Knowledge-Based Systems, Science of the Total Environment by Elsevier, each with two published papers. The majority of the corresponding authors are affiliated with institutions in China.
@@ 5.2.2 Datasets
AgroData-10 [art_28] https://www.aoml.noaa.gov/phod/argo
Carbon Monitor [art_360] https://carbonmonitor.org/
CCMP wind data [art_31], [art_40], [art_44] http://data.remss.com
China Environmental Monitoring Station PM2.5 [art_2], [art_8], [art_371] https://www.cnemc.cn/
China National Urban Air Quality [art_1], [art_5] https://github.com/Friger/GAGNN
Copernicus 3D thermohaline data [art_35] https://resources.marine.copernicus.eu/products
ECMWF East China Sea [art_45] https://github.com/Boltzxuann/GIPMN/tree/dataset
ECMWF South and East China Sea [art_32] http://apps.ecmwf.int/datasets/data/interim-full-daily
HadCRUT5 [art_370] https://doi.org/10.1029/2019JD032361 (source paper)
Jinan PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
KnowAir [art_6] https://github.com/shuowang-ai/PM2.5-GNN
Korean Peninsula sea data [art_33] https://www.nifs.go.kr/risa/main.risa
NOAA sea datasets [art_28], [art_36], [art_39], [art_46], [art_47], [art_365] https://www.ndbc.noaa.gov/
Norwegian offshore wind [art_29], [art_34] https://frost.met.no/index.html
Overbetuwe groundwater data [art_359] https://github.com/mlttac/GroundwaterFlowGNN
Taiwan air quality dataset [art_3] https://data.gov.tw/en
UrbanAir [art_6] https://doi.org/10.1016/j.chemosphere.2018.12.128 (source paper)
US EPA PM2.5 [art_1], [art_4] https://aqs.epa.gov/aqsweb/airdata/download_files.html
US Geological Survey [art_355] https://earthquake.usgs.gov/earthquakes/search/
Yangtze River PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
Six of the selected papers [art_28, art_36, art_39, art_46, art_47, art_365] use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S. Department of Commerce (US DOC). The US DOC website permits users to download daily, weekly and monthly mean optimum interpolation sea surface temperature (OISST) data from September 1981 for many geographical locations. In over half of the papers, exogenous variables are incorporated into the model. These include weather data (e.g., humidity, rainfall, pressure, temperature, wind) and, in the case of PM2.5 concentration forecasting, concentration of pollutants (like CO, NO2, O3, PM10 and SO2).
@@ 5.2.3 Types of models
Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNNs (28 out of 43 papers), followed by 9 attentional GNNs, and by 3 hybrid architectures. The description in 3 of the papers (Refs. [art_5, art_354, art_361]), however, does not allow us to specify the classification of the model within the proposed taxonomy. As for the graph structure, not all of the papers describe it with sufficient precision, and some of them are vague about the graph construction and its static or dynamic nature. In almost half of the papers, the edge weights are learned directly from the model, often using the graph attention mechanism. When the graph structure is defined manually, the most popular criterion used is the spatial distance between the sites ($d$), even in its exponential ($e^{-d}$) or inverse proportion ($1/d$) formulation, or similarity measures, such as cosine similarity and correlation of the time series. Finally, there are a few papers where more than one graph structure is fused together in order to capture more complex dynamics, as in Ref. [art_2]. Almost all the papers in this thematic group deal with multivariate time series, with the exception of Ref. [art_370], which studies a univariate time series of global average temperature. In this work, the authors treat each time step $t$ as a node $v_{t}$ in the graph, and the adjacency matrix is constructed in such a way that $v_{t}$ is influenced by $v_{t-1}$ and $v_{t+1}$.
@@
Dynamic Spatial-Temporal Graph Convolutional Recurrent Network (DSTGCRN) [art_360] https://github.com/PalaceTony/DSTGCRN
Dynamic spatiotemporal graph neural network (DST-GNN) [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
Group-aware graph neural network (GAGNN) [art_1] https://github.com/Friger/GAGNN
Gridded information propagation and mixing network (GIPMN) [art_45] https://github.com/Boltzxuann/GIPMN/tree/dataset
GNNCoder [art_356] https://doi.org/10.5281/zenodo.13358007
GroundwaterFlowGNN [art_359] https://github.com/mlttac/GroundwaterFlowGNN
Hierarchical graph recurrent network (HiGRN) [art_28] https://github.com/Neoyanghc/HiGRN
Physics-aware graph networks (PaGN) [art_365] https://github.com/eunviho/novel_pagn
Spatio-temporal FFTransformer (ST-FFTransformer) [art_29] https://github.com/LarsBentsen/FFTransformer
@@ 5.2.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_28], [art_33], [art_34], [art_38], [art_363], [art_370], [art_371]
Historical average (HA) [art_3], [art_6], [art_28], [art_38]
Naive [art_29], [art_34], [art_366]
Vector autoregression (VAR) [art_28], [art_34], [art_38], [art_360]
@@
Non-GNN machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_33], [art_46]
Combined FC-LSTM and convolution neural network (CFCC-LSTM) [CFCC-LSTM] [art_35], [art_39], [art_370]
DLinear [DLinear] [art_31], [art_364]
@@
Hybrid convolutional neural network and gated recurrent unit (CGRU) [CGRU] [art_31], [art_40], [art_44]
Convolutional neural network (CNN) [art_30], [art_46]
@@
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_4], [art_41], [art_368], [art_371]
Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_28], [art_41], [art_355]
Feed-forward neural network (FNN) [art_2], [art_7], [art_38], [art_360], [art_365], [art_370]
Gated recurrent unit encoder–decoder (GED) [GED] [art_36], [art_39], [art_47]
Gated recurrent unit (GRU) [art_2], [art_6], [art_7], [art_8], [art_38], [art_43], [art_46], [art_361], [art_364], [art_365], [art_367], [art_370]
Informer [Informer] [art_30], [art_47], [art_354], [art_364], [art_371]
Long-short term time series network (LSTNet) [LSTNet] [art_355], [art_373]
Predictive deep convolutional neural network (PDCNN) [PDCNN] [art_31], [art_40]
Random forest (RF) [art_6], [art_364], [art_370]
Support vector regression (SVR) [art_3], [art_6], [art_35], [art_36], [art_38], [art_360], [art_370], [art_371]
Recurrent neural network (RNN) [art_46], [art_365], [art_367]
Temporal convolutional network (TCN) [TCN] [art_6], [art_30], [art_355], [art_356], [art_368], [art_371], [art_373]
@@
Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_6], [art_31], [art_32], [art_35], [art_40], [art_44]
Transformer [attention] [art_30], [art_33]
Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [Wavelet-DBN-RF] [art_31], [art_40], [art_44]
Extreme gradient boosting (XGBoost) [art_1], [art_2], [art_5], [art_46], [art_364]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_6], [art_360], [art_365], [art_368]
@@
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_363], [art_368]
Deeper graph convolutional network (DeeperGCN) [DeeperGCN] [art_1], [art_5]
@@
Graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] [art_1], [art_2], [art_5], [art_6], [art_7], [art_354], [art_361]
Graph convolutional network (GCN) [art_38], [art_41], [art_357], [art_367]
Graph WaveNet (GWN) [GWN] [art_1], [art_6], [art_44], [art_47], [art_363], [art_368]
Hierarchical graph convolution network (HGCN) [HGCN] [art_1], [art_28]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_6], [art_32], [art_44], [art_355], [art_360], [art_373]
Multi long-short term memory network (Multi-LSTM) [Multi-LSTM] [art_31], [art_40], [art_44]
Superposition graph neural network (SGNN) [art_24] [art_31], [art_40]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_3], [art_6], [art_28], [art_35], [art_47], [art_363], [art_368], [art_371]
Spatio-temporal correlation graph neural network (STGN) [art_40] [art_31], [art_44]
Spatio-Temporal U-Network (ST-UNet) [ST-UNet] [art_1], [art_5]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_368], [art_371]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_6], [art_30], [art_35]
@@
The majority of the benchmarks used in this thematic group fall within the category of non-GNN machine learning methods, although there are several GNN benchmarks as well. The two most common benchmark models are LSTMs and GRUs. These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and are considered a reliable benchmark in the field. As for GNN models, the most common benckmarks are the graph convolutional neural network and long-short term memory newtork (GC-LSTM) [GC-LSTM] (which appears only in quality air forecasting problems), and the spatio-temporal graph convolutional network (STGCN) [STGCN]. The GC-LSTM network is a model that integrates graph convolutional networks and long-short-term memory networks. It was proposed by Qi et al. in 2019 to forecast the spatio-temporal variation of PM2.5 concentration. In the original paper, the GC-LSTM model works with a pre-defined graph based on the geographical distance between air quality stations. The input layer of the GC-LSTM model receives historical air quality observations, meteorological variables, spatial predictors, and temporal predictors. This information is processed through a graph convolution operation to extract spatial features. After that, the temporal feature extraction is performed using a LSTM network, which takes as input the graph convolutional features concatenated with the original signals. The STGCN model, on its part, was proposed in the traffic domain by Yu et al. in 2018, and works with pre-defined adjacency matrices computed based on the distances among stations in the traffic network. The STGCN model is composed of two spatio-temporal convolutional blocks and a fully-connected output layer. Each convolutional block contains two temporal gated convolution layers and one spatial graph convolution layer in the middle. The code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. Finally, it is important to point out that the so-called multivariate time series forecasting with graph neural network (MTGNN) model proposed by Wu et al. in 2020 in Ref. [MTGNN] is often reported in the papers using it as the second-best performing model, thus it probably deserves special attention and further investigation in future research.

### v4

@@ 5.1.5 Results
In general, the authors of the papers in this group claim that their proposed GNN model outperforms the benchmarks. However, in some cases, it turns out that basic statistical methods may achieve good results as well [319]. The most common error metrics used are the MAE, the RMSE, and the Normalized Root mean squared Error (NRMSE). Some papers calculate these error metrics across different seasons to evaluate the model’s performance under different conditions. An example is provided in Ref. [377], where the authors also use the GNNExplainer [455] to assess the influence of different variables.
It is not possible to compare the accuracy of the models included in this group, because no common benchmark datasets were used. The only two papers that study the same dataset are written by the same authors (Refs. [350, 351]), and their most recent paper includes a comparison with the models proposed in the first one.
@@ 5.2 Environment
The thematic group “Environment” includes a relevant number of selected papers (43 out of 366). This suggests a considerable research interest on the application of GNNs in this field. However, this group is highly fragmented, as it includes a large number of subdomains and applications, all related to the study of environmental data.
@@ 5.2.1 Overview
The group of papers related to environmental topics is broad, and includes papers on disciplines such as physics, climatology, oceanography, and atmospheric science. The breadth of the group is reflected in the number of case studies, which cover air quality prediction (in terms of PM2.5 concentration), sea temperature, wind speed, and other applications. Forecasting these quantities can be useful for several reasons. For instance, predicting wind speed or rainfall can assist in anticipating some extreme weather phenomena, and is also related to the prediction of green energy outputs. For example, in Refs. [11] and [12] the authors estimate the wind power forecasting error obtained by using wind speed forecasting models. As another example, predicting sea surface temperature can be useful for weather forecasting, fishing directions, and disaster warnings [437]. Moreover, monitoring and forecasting environmental data are crucial for evaluating the impact of human activities on the environment and tracking the progression of climate change.
@@
An overview of the selected papers reveals that there has been a rapid increase in interest on this theme among the GNN research community too, and more than 35 of the selected papers were published from 2023. The most investigated data are related to sea temperature (Refs. [437, 185, 304, 98, 392, 361, 345, 42]), wind speed (Refs. [11, 326, 12, 102, 97, 73, 269]), and PM2.5 concentration forecasting (Refs. [267, 357, 223, 293, 183, 310, 248, 325, 426]). Almost every selected paper is published in a different journal, with the exception of Energy, Engineering Applications of Artificial Intelligence, Journal of Cleaner Production, Knowledge-Based Systems, Science of the Total Environment by Elsevier, each with two published papers. As for conference papers, the three “Environment” papers were presented at KDD, IJCAI and CIKM.
@@ 5.2.2 Datasets
AgroData-10 [437] https://www.aoml.noaa.gov/phod/argo
Carbon monitor [106] https://carbonmonitor.org/
CCMP wind data [326], [102], [97] http://data.remss.com
China environmental monitoring station PM2.5 [267], [310], [426] https://www.cnemc.cn/
China national urban air quality [41], [72] https://github.com/Friger/GAGNN
Copernicus 3D thermohaline data [304] https://resources.marine.copernicus.eu/products
Delaware river basin [42] https://waterdata.usgs.gov/nwis
ECMWF east China sea [311] https://github.com/Boltzxuann/GIPMN/tree/dataset
ECMWF south and east China sea [502] http://apps.ecmwf.int/datasets/data/interim-full-daily
HadCRUT5 [403] https://crudata.uea.ac.uk/cru/data/temperature/
HUC12 [5] https://hawqs.tamu.edu/
Jinan PM2.5 [248] https://github.com/liaohaibing/MGST_GNN/tree/master
KnowAir [293] https://github.com/shuowang-ai/PM2.5-GNN
Korean peninsula sea data [185] https://www.nifs.go.kr/risa/main.risa
NARR weather conditions [133] https://psl.noaa.gov/data/gridded/data.narr.html
NOAA sea datasets [437], [98], [392], [73], [345], [336] https://www.ndbc.noaa.gov/
Norwegian offshore wind [11], [12] https://frost.met.no/index.html
Overbetuwe groundwater data [366] https://github.com/mlttac/GroundwaterFlowGNN
Taiwan air quality dataset [357] https://data.gov.tw/en
UrbanAir [293] https://doi.org/10.1016/j.chemosphere.2018.12.128 (source paper)
US EPA PM2.5 [41], [223] https://aqs.epa.gov/aqsweb/airdata/download_files.html
WeatherBench [133] https://github.com/pangeo-data/WeatherBench
Yangtze river PM2.5 [248] https://github.com/liaohaibing/MGST_GNN/tree/master
Six of the selected papers [437, 98, 392, 73, 345, 336] use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S. Department of Commerce (US DOC). The US DOC website permits users to download daily, weekly and monthly mean optimum interpolation sea surface temperature (OISST) data from September 1981 for many geographical locations. In over half of the papers, exogenous variables are incorporated into the model. These include weather data (e.g., humidity, rainfall, pressure, temperature, wind) and, in the case of PM2.5 concentration forecasting, concentration of pollutants (like CO, NO2, O3, PM10 and SO2).
@@ 5.2.3 Types of models
Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNNs (28 out of 43 papers), followed by 8 attentional GNNs, and by 3 hybrid architectures. The description in 4 of the papers (Refs. [72, 248, 325, 42]), however, does not allow us to specify the classification of the model within the proposed taxonomy. As for the graph structure, not all of the papers describe it with sufficient precision, and some of them are vague about the graph construction and its static or dynamic nature. In almost half of the papers, the edge weights are learned directly from the model, often using the graph attention mechanism. When the graph structure is defined manually, the most popular criterion used is the spatial distance between the sites ($d$), even in its exponential ($e^{-d}$) or inverse proportion ($1/d$) formulation, or similarity measures, such as cosine similarity and correlation of the time series. Finally, there are a few papers where more than one graph structure is fused together in order to capture more complex dynamics, as in Ref. [267]. Almost all the papers in this thematic group deal with multivariate time series, with the exception of Ref. [403], which studies a univariate time series of global average temperature. In this work, the authors treat each time step $t$ as a node $v_{t}$ in the graph, and the adjacency matrix is constructed in such a way that $v_{t}$ is influenced by $v_{t-1}$ and $v_{t+1}$.
@@
Domain inspired temporal graph convolution neural network [5] https://github.com/IBM/tgcn-soil-moisture
Dynamic spatial-temporal graph convolutional recurrent network (DSTGCRN) [106] https://github.com/PalaceTony/DSTGCRN
Dynamic spatiotemporal graph neural network (DST-GNN) [248] https://github.com/liaohaibing/MGST_GNN/tree/master
Group-aware graph neural network (GAGNN) [41] https://github.com/Friger/GAGNN
Gridded information propagation and mixing network (GIPMN) [311] https://github.com/Boltzxuann/GIPMN/tree/dataset
GroundwaterFlowGNN [366] https://github.com/mlttac/GroundwaterFlowGNN
Hierarchical graph recurrent network (HiGRN) [437] https://github.com/Neoyanghc/HiGRN
Physics-aware graph networks (PaGN) [336] https://github.com/eunviho/novel_pagn
Spatio-temporal FFTransformer (ST-FFTransformer) [11] https://github.com/LarsBentsen/FFTransformer
@@ 5.2.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [437], [185], [12], [257], [416], [403], [426], [133]
Historical average (HA) [357], [293], [437], [257], [133]
Naive [11], [12], [269]
Vector autoregression (VAR) [437], [12], [257], [106]
@@
Non-GNN machine learning methods Bidirectional long-short term memory network (BiLSTM) [109] [185], [73]
Combined FC-LSTM and convolution neural network (CFCC-LSTM) [442] [304], [392], [403]
DLinear [469] [326], [417]
@@
Hybrid convolutional neural network and gated recurrent unit (CGRU) [433] [326], [102], [97]
Convolutional neural network (CNN) [210], [73]
@@
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [223], [361], [45], [426]
Convolutional long-short term memory network (ConvLSTM) [346] [437], [361], [229]
Feed-forward neural network (FNN) [267], [183], [257], [106], [336], [403]
Gated recurrent unit encoder–decoder (GED) [430] [98], [392], [345]
Gated recurrent unit (GRU) [267], [293], [183], [310], [257], [10], [73], [325], [417], [336], [403], [133]
Informer [520] [210], [345], [248], [417], [426], [133]
Long-short term memory network (LSTM) [41], [267], [357], [223], [72], [293], [183], [310], [437], [210], [502], [185], [304], [98], [257], [392], [10], [73], [345], [184], [325], [417], [336], [45], [403], [426], [17], [42], [5], [133]
Long-short term time series network (LSTNet) [202] [229], [17]
Predictive deep convolutional neural network (PDCNN) [525] [326], [102]
Random forest (RF) [293], [417], [403]
Recurrent neural network (RNN) [73], [336]
Support vector regression (SVR) [357], [293], [304], [98], [257], [106], [403], [426]
Temporal convolutional network (TCN) [9] [293], [210], [229], [45], [426], [17]
@@
Temporal pattern attention long-short term memory network (TPA-LSTM) [347] [293], [326], [502], [304], [102], [97]
Transformer [375] [210], [185]
Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [164] [326], [102], [97]
Extreme gradient boosting (XGBoost) [41], [267], [72], [73], [417]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [293], [106], [336], [45]
@@
Attention-based spatial-temporal graph convolutional network (ASTGCN) [116] [416], [45], [133]
Diffusion convolutional recurrent neural network (DCRNN) [237] [416], [133]
@@
Graph convolutional neural network and long-short term memory (GC-LSTM) [316] [41], [267], [72], [293], [183], [248], [325]
Graph convolutional network (GCN) [257], [361], [184]
Graph WaveNet (GWN) [425] [41], [293], [97], [345], [416], [45]
Hierarchical graph convolution network (HGCN) [115] [41], [437]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [424] [293], [502], [97], [229], [106], [17], [133]
Multi long-short term memory network (Multi-LSTM) [103] [326], [102], [97]
Superposition graph neural network (SGNN) [460] [326], [102]
Spatio-temporal graph convolutional network (STGCN) [457] [357], [293], [437], [304], [345], [416], [45], [426], [133]
Spatio-temporal correlation graph neural network (STGN) [102] [326], [97]
Spatio-temporal synchronous graph convolutional network (STSGCN) [353] [45], [426]
Temporal graph convolutional network (T-GCN) [508] [293], [210], [304], [5], [133]
@@
The majority of the benchmarks used in this thematic group fall within the category of non-GNN machine learning methods, although there are several GNN benchmarks as well. The two most common benchmark models are LSTMs and GRUs. These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and are considered a reliable benchmark in the field. As for GNN models, the most common benchmarks are the graph convolutional neural network and long-short term memory network (GC-LSTM) [316] (which appears only in quality air forecasting problems), and the spatio-temporal graph convolutional network (STGCN) [457]. The GC-LSTM network is a model that integrates graph convolutional networks and long-short-term memory networks. It was proposed by Qi et al. in 2019 to forecast the spatio-temporal variation of PM2.5 concentration. In the original paper, the GC-LSTM model works with a pre-defined graph based on the geographical distance between air quality stations. The input layer of the GC-LSTM model receives historical air quality observations, meteorological variables, spatial predictors, and temporal predictors. This information is processed through a graph convolution operation to extract spatial features. After that, the temporal feature extraction is performed using a LSTM network, which takes as input the graph convolutional features concatenated with the original signals. The STGCN model, on its part, was proposed in the traffic domain by Yu et al. in 2018, and works with pre-defined adjacency matrices computed based on the distances among stations in the traffic network. The STGCN model is composed of two spatio-temporal convolutional blocks and a fully-connected output layer. Each convolutional block contains two temporal gated convolution layers and one spatial graph convolution layer in the middle. The code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. Finally, it is important to point out that the so-called multivariate time series forecasting with graph neural network (MTGNN) model proposed by Wu et al. in 2020 in Ref. [424] is often reported in the papers using it as the second-best performing model, thus it probably deserves special attention and further investigation in future research.